OpenStudy (anonymous):

Find the exact value of the radical expression in simplest forms (Sqrt)16y3-(sqrt)y3

OpenStudy (campbell_st):

if \[\sqrt{16y^3} = 4y^(3/2) \] \[\sqrt{y^3} = y^{3/2}\] so its \[(4 -1)y^{3/2} = 3y^{3/2}\]

OpenStudy (anonymous):

Ok thanks